An individual has $50,000 invested in a stock with a beta of 0.5 and another $70,000 invested in a stock with a beta of 1.4. If these are the only two investments in her portfolio, what is her portfolio's beta? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1
Total Portfolio value = Value of Stock 1 + Value of Stock 2
=50000+70000
=120000
Weight of Stock 1 = Value of Stock 1/Total Portfolio Value
= 50000/120000
=0.4167
Weight of Stock 2 = Value of Stock 2/Total Portfolio Value
= 70000/120000
=0.5833
Beta of Portfolio = Weight of Stock 1*Beta of Stock 1+Weight of Stock 2*Beta of Stock 2
Beta of Portfolio = 0.5*0.4167+1.4*0.5833
Beta of Portfolio = 1.03